OpenStudy (anonymous):

Nicole’s Expenses rent x% car y% other 19% food 18% education 20% Based on the graph above, if Nicole spends $1000 each month, how much does she spend on her car and rent?

OpenStudy (anonymous):



OpenStudy (anonymous):

The graph roughly looks like that.

OpenStudy (anonymous):

we know that she only has 100% to use in percentage wise so: 100%=20%+18%+19%+x%+y% we also know that the total money is: $1000=18%+19%+20%+y%+x% i would simplify each of the percentages on the second equation as the first part: try just that and post your answer here.

OpenStudy (anonymous):

$1000= 57% +x%+y%

OpenStudy (anonymous):

ok now how much is 57% of $1000?

OpenStudy (anonymous):

570

OpenStudy (anonymous):

ok try simplifying the top equations.

OpenStudy (anonymous):

so you get just y on one side

OpenStudy (anonymous):

430-x%=y%

OpenStudy (anonymous):

not quite do you see where the mistake is taking a second look at it?

OpenStudy (anonymous):

The percent?

OpenStudy (anonymous):

the number yeah how would you fix it?

OpenStudy (anonymous):

Multiply it to the other side?

OpenStudy (anonymous):

Actually you would divide it...I think...

OpenStudy (anonymous):

never mind i see what you did there wasnt a mistake just a confusion. i had said the first problem but you simplified the second sorry just got confused there ok so you were right. ok so lets take that one and plug it into the other equation

OpenStudy (anonymous):

100= 430 -x% = y%

OpenStudy (anonymous):

ok sort of on the right track but what you want to do is keep everything else that was there so: 100%=20%+18%+19%+x%+y% and then plug in the $430 -x% = y% so it would look like this: 100%=20%+18%+19%+x%+$430 -x%

OpenStudy (anonymous):

100%=57%+x%+430-x%

OpenStudy (anonymous):

100%=57%+430

OpenStudy (anonymous):

43%=430

OpenStudy (anonymous):

=430

OpenStudy (anonymous):

so 430 is the total for both rent and car the only way to get simpler would be to have more info 430=car+rent not sure if that is what they want

OpenStudy (anonymous):

430 is one of the answers